OpenStudy (anonymous):

How do you integrate this?? (9x^2)/(1-x^3)^0.5

hartnn (hartnn):

put 1-x^3 = u du =.. ?

hartnn (hartnn):

did u try that ? u = 1-x^3 @mbhatti_95

OpenStudy (anonymous):

yes I did but i keep getting stuck... can you please help me with the solution?

hartnn (hartnn):

sure. what did u get du as ... ? or which step are u stuck ?

OpenStudy (anonymous):

du = -3x^2 dx i am stuck after the integration when you have to give answer in terms of x... :S

hartnn (hartnn):

so u could integrate properly, right ? so u might have got -2/3 \(\sqrt u\)+c now again put u=1-x^3 -2/3 \(\sqrt{1-x^3} +c \)

OpenStudy (anonymous):

i keep getting -27( (1-x^3) - 2/3((1 - x^3)^3/2))

hartnn (hartnn):

ok, after substitution did u get \(\large \int \frac{-1}{3\sqrt u}du\) ?

OpenStudy (anonymous):

allright pellet... I realised my mistake... my bad sorry.. got it thanks a ton!!

hartnn (hartnn):

good. welcome ^_^